Chapter 2, Problem 2/108 The sphere Ptravels in a straight line with speed v-9.7 m/s. For the instant depicted, determine the corresponding values of r and θ as measured relative to the fixed Oxy coordinate system. 5.0 m 4.0 m 0 Answers: m/s rad/s
Answer #1

Given,

v = 9.7 m/s

theta = tan^-1(4/5) = 38.66 deg

angle = 90 - theta - 31 = 20.34 deg

vr = -v cos(angle) = -9.7 x cos20.34 = -9.1 m/s

v(theta) = -9.7 sin20.34 = -3.37 m/s

theta = v(theta)/sqrt (4^2 + 5^2) = -3.37/6.4 = - 0.527 rad/s